OpenStudy (anonymous):

You deposit $1,000 for 4 years at an interest rate of 2%. If the interest is compounded quarterly, how much money would you have after the four years?

OpenStudy (anonymous):

\[1000\times \left(1+\frac{.02}{4}\right)^{2\times 4}\] and a calculator

OpenStudy (anonymous):

I was checking my ans. I got $1083.10 the I also compounded with annual int and got $1082.40